### v1

# Dual-Branch Convolutional Framework for Spatial and Frequency-Based Image Forgery Detection

Abstract
With a very rapid increase in deepfakes and digital image forgeries, ensuring the authenticity of images is becoming increasingly challenging. This report introduces a forgery detection framework that combines spatial and frequency-based features for detecting forgeries. We propose a dual branch convolution neural network that operates on features extracted from spatial and frequency domains. Features from both branches are fused and compared within a Siamese network, yielding 64 dimensional embeddings for classification. When benchmarked on CASIA 2.0 dataset, our method achieves an accuracy of 77.9%, outperforming traditional statistical methods. Despite its relatively weaker performance compared to larger, more complex forgery detection pipelines, our approach balances computational complexity and detection reliability, making it ready for practical deployment. It provides a strong methodology for forensic scrutiny of digital images. In a broader sense, it advances the state of the art in visual forensics, addressing an urgent requirement in media verification, law enforcement and digital content reliability.
1 Introduction and Background
Digital images are a significant part of communication, sharing information, and decision-making in a variety of fields and industries, including journalism, social media, surveillance, and forensics. Unfortunately, advances in image editing applications and the availability of AI-centric methods of generation have exaggerated the ease with which imagery can be manipulated to mislead its viewers and compromise honesty. It is important to contextualize this within a proper definition of image tampering, which means any alteration or distortion added to an original digital image to alter the messages being conveyed or develop a false narrative. There is a spectrum of common manipulation techniques that may include simple edits of cropping, resizing, and color change, simple but advanced methods like splicing (combination of content from different images together), copy–move (copying over a portion of the image), inpainting (filling in a missing area of an image), and most recently, the creation of deepfakes using generative modeling.
The history of digital image forgeries began sometime after photography came into existence with darkroom techniques being used to delete or add elements. First, digital editing programs such as Adobe Photoshop normalized the use of digital imaging and gave a world of possibilities to anyone willing to whether for positive or negative intent. The shift from manual image editing to fully automated, AI-assisted image editing methods means that conceptualizing and implementing image tampering is significantly less laborious and is now less detectable. The rapid growth of deep learning, GANs, and diffusion models have brought us to a point in the timeline of forgeries, where real synthetic images can be generated quickly, that bear a highly realistic look and semantically accurate effect and where traditional detection methods are practically useless.
In this scenario, one of the main obstacles in digital forensics has been to authenticate and validate digital content, particularly digital images. The exponentially increasing volume and availability of altered media on online social platforms has created serious issues with verifying visual evidence. The difficulty is not simply formulating a way to identify obvious alterations, but also detecting subtle, context-sensitive alterations that could fool conventional statistical methods of detection. The evolution of image creation technologies presents an equal and parallel evolution in the means and tools to provide verification.
The purpose of this paper is to examine the state of the art in detection of tampering of digital images, looking at both traditional as well as deep-learning methods. The objectives of this study are three-fold: first, to review and classify approaches in the literature; second, to evaluate the feasibility and limitations of various techniques toward contemporary image manipulation methods; and third, to reveal existing voids and open challenges that need to be filled to produce effective, scalable, and trustworthy detection methods. Overall, this discussion aims to be helpful in the continued efforts to preserve trust and integrity in digital imagery in a time where visual synthesis is unprecedented.
2 Literature Review
The detection of digital image tampering has been a long-standing area of research in digital forensics [1], gaining prominence as editing tools have become more sophisticated and accessible. The early methods in this field relied largely on handcrafted statistical and signal-processing techniques [2]. These methods aimed to expose inconsistencies introduced during the editing process, such as abrupt changes in lighting, irregular noise patterns, or compression artifacts. Techniques like error level analysis (ELA) [3], discrete cosine transform (DCT) [4] coefficient inspection, and sensor pattern noise analysis [5] were developed to identify these inconsistently manipulated areas. They laid the foundation for automated detection pipelines and demonstrated the feasibility of identifying image tampering using statistical measures. However, their effectiveness was often limited to specific types of edits, making them struggle to generalize across diverse image sources and varying post-processing conditions. [6]
As the field progressed, tampering detection methods diverged into two broad categories: passive (or blind) and active approaches. Passive methods operate exclusively on the received images, relying on statistical, structural, or perceptual inconsistencies to locate manipulated areas [7]. These approaches gained widespread adoption due to their ability to work with any digital image, regardless of its source, making them ideal for forensics applications across platforms and media. In contrast, active methods embed digital watermarks, hash codes, or unique identifiers at the time of image capture [8]. These embedded marks enable verification and tamper detection with a high degree of reliability. However, active methods require access to or control of the original camera or editing process, making them impractical for situations where images are circulated online or captured by arbitrary devices. As a result, passive methods have emerged as the dominant approach for digital tampering detection.
With advances in machine learning and deep learning, tampering detection has entered a new era. Convolutional neural networks (CNNs) [9, 10] and, more recently, transformers [11] have demonstrated remarkable capabilities for learning hierarchical, context-aware features that expose inconsistencies introduced during editing. These deep learning methods have proven highly effective for challenging tampering scenarios, including splicing, copy-move, inpainting, and deepfake generation. Models such as multi-task fully convolutional networks have demonstrated strong performance in locating tampered areas by focusing on noise inconsistencies and spatial artifacts. Meanwhile, noiseprint-based approaches [12] have leveraged deep feature embeddings to isolate camera or sensor characteristics, making it possible to pinpoint even highly sophisticated manipulations. Newer methods utilize attention mechanisms, self-supervised learning, and multi-scale feature aggregation to further improve detection and generalization across diverse datasets and editing techniques [13, 14].
Despite significant advances, the field still faces a range of critical challenges. The emergence of highly realistic, AI-generated images has pushed the boundaries of traditional detection pipelines [15], making it increasingly difficult to distinguish between authentic and manipulated media. The scarcity of large-scale, high-quality, and well-annotated tampering datasets also limits the ability of deep learning models to learn robust and generalizable representations. In addition, the vulnerability of deep detection methods to adversarial attacks presents a serious concern, as carefully crafted perturbations can fool state-of-the-art detection systems [16]. Another challenge lies in extending detection methods to multi-modal settings, such as verifying the consistency between images and associated text, or detecting inconsistencies across frames in a video [17, 18, 19].
For future work, it will be important to address these gaps by exploring resilient, multi-modal methods for detection that will have reliable performance across a variety of operating conditions. It will be essential to develop explainable AI approaches for tampering detection so that investigators will have insight into the evidence generated by deep learning-based detectors and trust in automated evaluations. In addition, expanding tampering detection work to also include video and cross-modal data - and improving on anomaly detection - will better prepare the field for the various threats posed by ever-more advanced and context-oriented manipulation techniques.
3 Image Tampering Taxonomy
Digital image tampering can encompass a diversity of techniques that alter visual content for deceptive or misleading purposes. To understand the landscape of these techniques will be valuable to select and develop appropriate detection techniques. In this section, we will outline the common classes of image tampering, their features, and the types of forensic difficulties.
3.1 Copy–Move Forgeries
Copy–move forgery is one of the oldest and most common forms of image tampering in the digital domain. In a copy–move forgery, a part of an image is copied and pasted in some other location of the same image. Many times the objective is to hide or duplicate something that has an intention to make the original scene either look different or not part of the original scene. Copy–move forgeries are hard to detect, especially when the image has been subject to post-processing steps like geometric deformation, rotation, or change in size. The copied image part can be changed in scale, obtained by flipping the image piece, or doing a slight affine deformation so that it fits seamlessly into other parts of the scene. These manipulations can prevent detection, by statistical or block matching methods, of the originally copied image part. From a forensic perspective, it is important to detect copy–move forgery as a forensic analysis to establish whether an image that is being presented to confirm events, such as published images, judicial images, and surveillance images, accounts for the integrity of the original image.
Figure 1: (a) Real image (b) Tampered image
3.2 Splicing and Composition
Splicing is the removal of parts of one or more source images and combining them into the target image to create a new, realistic scene. Advanced splicing can create images that will evade legacy detection methodologies, and manipulated images will usually introduce some inconsistency into the image. For instance, differences in lighting, noise, or color balance may exist between sections. In some cases, there could be evidence at the boundaries of the spliced regions that could be revealed with edge detection or error level analysis methods. In the current environment of deep generative models and AI support in image editing, splicing has become highly realistic and has become one of the more challenging forms of image forgery to demonstrate in digital image forensics.
Figure 2: Generation of a forged image using splicing
3.3 Retouching and Enhancement
When we edit and enhance images, we are operating at the pixel or object level with highly localized changes, ranging from minor tweaks of color, brightness, and contrast to more substantial changes such as insertion, replacement, or removal of objects. We are able to conceal or emphasize aspects of images to change the perceived context of the content. Inpainting, airbrushing, and ”automatically filling in” objects are commonplace techniques with general image editing packages, and their widely-accessible nature facilitates widespread forgery in media and advertising. The problem means processing and processing are very localized - in fact, the rapid growth of ease of automated edits and deep learning fueled further challenges predicting and identifying original images. It also requires attention to more sophisticated statistical and deep learning based approaches to identify subtle variations in texture, lighting, and noise.
Figure 3: (a) Real image (b) Retouched image
3.4 Geometric Transformations
The category of geometric transformations refers to a variety of manipulations we can make to alter the spatial properties of an image, including scaling, rotation, cropping, warping, and perspective distortion. In fact, editing an image can change not only the spatial relationships between objects but also distort the fundamental structure of the pixels in the image. Common artifacts of these types of manipulations include aliasing effects caused by resampling, discontinuities caused by an arbitrary change to an image’s boundary, and repeating patterns which indicate repeated resizing or compression. In many cases lossless compression adds additional noise and makes it even more challenging to detect inconsistencies in manipulated areas. To assess the truth of an image, it is crucial to identify whatever geometric transformation has occurred, particularly if the image has implications in surveillance, intelligence, or digital evidence.
In conclusion, an understanding of the nature, use and forensic attributes of copy–move, splicing, retouching and geometric transformations provides a solid basis for addressing the increasingly complex issues arising from image manipulation. Developments in artificial intelligence and deep learning, in particular, have made it more difficult to distinguish between unaltered and manipulated content, making solid, multi-modal detection approaches an urgent need for research and development.
4 Proposed Methodology: Hybrid Noise-Spatial-Frequency Siamese Network
4.1 Overview of the Proposed Approach
The proposed methodology presents a unique hybrid implementation that simultaneously leverages spatial and frequency domain feature extraction method within a Siamese network architecture for robust detection of image splicing forgeries. The new feature extraction method is based on a joint deep learning approach that differentiates itself from earlier models by using the ability to integrate features from both spatio and frequency domain feature spaces, resulting in better discriminative capabilities. The two-pathway network runs all proposed key patches of an image through one of two Siamese networks to learn the characteristics of an original and manipulated part of an image, within an end-to-end learning paradigm applied to the challenging splicing forgery in the CASIA 2.0 dataset. Overall, this method showed that a unique hybrid method for detecting image splicing was successful in learning the features needed to achieve such robust detection performance.
4.2 Noise Feature Extraction
The Noise Extraction Module serves to expose high-frequency artifacts, frequently indicative of manipulations, including splicing, or subtle image inferences. As a PyTorch module, it consists of a learnable set of high-pass filters, initialized from well-known operators - the Laplacian and Sobel filters (along both X and Y directions). These filters modify pixel intensity discontinuities to represent edges and changes at these fine levels. Together with fixed filters, the module employs randomly initialized filters that are learnt through training, allowing the network to adaptively learn noise for patterns that suggest tampering. The input, typically upper-channel RGB, is converted to a grey-scale representation for frequency limiting convenience. The grey scale representation image passes through a set of high-pass filters via 2D convolution, gaining noise maps that are skewing from the not-trivial object and its surroundings, in which they highlight non-natural high-frequency components. After convolving, a batch normalization layer stabilizes the output, and the hyperbolic tangent activation force the noise responses into a fixed range, aiding downstream learning initiatives of low-frequency histograms. The noise extraction module is crucial to indirectly increasing the model’s attention to rely on low-frequency elements of various processes, to assist the actual identification of inconspicuous cues identified for image-processing oblivion; it is important to highlight its value in the overall forensic division.
Figure 4: Noise Extraction Module Architecture
4.3 Spatial Feature Extraction
The proposed method generally uses a variety of spatial features and statistical measures that can indicate the different artifacts of an image splice. One of the methods utilized is a Local Binary Pattern and other variations. LBPs detect discrepancies in micro-textures that often arise when stitching together two different regions of an image. While humans often ignore the unnoticeable differences in textures, it is immutable based on LBP based descriptor analysis. Next, statistical measures are established based on edges that develop at probable boundaries of a splice to determine the correct structure of the image (i.e., continuity of edges across a splice). The edge profile follows both leap or gradual changes that either represent obvious signs of modification or a lack of seamlessness as might be established in detectable poor blending when different images are poorly blended together. This change could be seen with uneven lighting or contrasting colors. Along with edges, gradients of pixel intensities are evaluated in a similar approach to analyze the smoothness and directionality of the respective intensity transition. In pictures with altered sections, there could exist unnatural transitions in the gradual transition of pixels (e.g., aspects of the splice). Lastly, decomposing and analyzing the variance across patches will take advantage of differences in noise from signal across images. Since each of the patches involved in the splice comes from a likely different source or history of compression, then manipulating the different noise across each area should yield evidence of some degree of forgery.
Figure 5: Spatial Feature Extraction Architecture
4.4 Frequency Domain Feature Analysis
The frequency domain approach is to identify inconsistencies that the compression artifacts and edit have introduced by examining the underlying frequency properties of the image where documents will have an entirely different frequency distribution of content. Part of this process involves looking at the index of the Discrete Cosine Transform (DCT) coefficients and the distribution which can indicate statistical outliers that often identify areas that are spliced together. Each region in a spliced image will alter the normal statistics of DCT coefficients, so any reasonable deviation is a potential sign that tampering has occurred. In addition to this, an analysis of JPEG compression artifacts is carried out because consistently edited images often create artifacts along the block boundaries as a result of JPEG compression, because each region has JPEG compression and associated quantization tables based on each region’s respective pixel frequency, edited regions would lead the user to conclusions about the image that would not be correct and could yet, be entirely overlooked if analysed in the spatial domain. The approach also uses Fourier analysis to detect periodicity or disturbances in periodicity due to edits. For example, splicing may create an irregularity and outers will also be evident by altered periodicities. Wavelet decomposition can be performed to decompose the image into multiple levels of resolution. The ”residual” or high-frequency detail can also be isolated and analyzed across different scales and their noise characteristics with respect to splicing. Lastly, power spectral density (PSD) can be calculated and analysed across different regions of the image.
4.5 Feature Fusion Strategy
In order to create a robust and complete feature space for splicing detection, the method presented concerns full spatial and frequency-domain feature space with multi-step fusion. The first step is concatenation of features, whereby features obtained from the analysis on spatial and frequency features have become a single feature vector. It will help our system to use the strengths of each analysis, since both the texture inconsistencies of splicing and the frequency inconsistencies of splicing have been combined into a single feature vector. However, these feature levels will never completely correspond to each other, and we will notice that many of the elements will include both redundant or non-informative aspects or, collectively, they will make up high dimensions. To manage high dimensions we can, for instance, use Principal Component Analysis (PCA), or bottleneck layers in the set of auto encoders. These methods will condense the material we need while reducing any computational requirements, and should also enhance both the efficacy and efficiency of the subsequent stages of the detection pipeline. Following this step of condensation, both normalization, and in some instances scaling, are required to engage with the variation in features across sub-domains and categories. In step 2; we do not want a situation where one of our feature types overly dominates the input range due to the difference in scale, and considerations in the context of the input space are ongoingly engaged. Finally, feature selection is shared at this stage, and in all cases it is beneficial to pitch out features that do not give us discriminating power.
4.6 Siamese Network Architecture
The detection mechanism is based on a Siamese network type that is especially appropriate for pairwise comparison problems such as splicing detection. A Siamese network is made up of two parallel identical but separate subnetworks. Each subnetwork processes a patch of an image. The two branches have shared weights, ensuring that both patches are encoded in some shared, comparable representation space. Furthermore, the two identical branches enable consistent, meaningful comparisons regardless of the information they contain. A critical component of the learning framework will involve distance metric learning where a contrastive loss function can be leveraged to train the distance between the feature embeddings produced by the network. The aim is to find embedding pairs that are similar (genuine patches) have a minimized distance and have a maximized distance for dissimilar pairs (spliced patches versus genuine). The aim of these two learning conventions is for the network to find a collection of discriminative features that can differentiate spliced content from what is real. Similar to shared weight parameters encouraging consistent feature learning across the two subnetworks; however, it also encourages generalization, in addition to limiting the likelihood of overfitting. The architecture has also been developed and optimized to balance computing loads with detection capabilities by adjusting the depth and configuration of the layers in the networks.
4.7 Training and Optimization Protocol
The proposed technique was trained and validated with the benchmark CASIA 2.0 dataset, which is a standard in image splicing detection. To maximize the generalizability of the model across all types of manipulation, a full data preprocessing and augmentation pipeline was undertaken. The augmentation involved random cropping, horizontal flipping, color perturbations, and scaling which mimicked the various realistic forms of splicing. Additionally, a patch extraction scheme ensured balanced sampling from both legitimate and spliced regions, allowing the network to learn discriminative features and recognising splicing related dissimilar features. Balancing the sampling from both regions was significant so as not to bias the learning towards one class and promote robust detections across both classes. Learning was guided by hyperparameter tuning with respect to learning rate, batch size, and network depth to optimise the a stable ’convergence’ and performance. In addition to meticulously optimising parameters, the K-fold cross validation scheme was adopted to further measure the model for reliability and not overfitting. The K-Fold cross validation provided a comprehensive report of the models detection performance by changing validation sets across those subsets in the datasets. Collectively these strategies managed to maintain a stringent training environment and ensure performance was consistent across splicing.
4.8 Comparative Baseline Methods
To evaluate the efficacy of the proposed approach, a range of traditional and state-of-the-art methods are implemented as baselines:
  • Statistical Analysis Approaches: Techniques relying on pixel correlation, color distribution examination, and noise inconsistencies.
  • Traditional Frequency Domain Methods: Approaches based on DCT coefficient examination, JPEG artifact detection, and Fourier transform periodicity detection.
  • Conventional Machine Learning Approaches: Support Vector Machines (SVM) and Random Forest classifiers utilizing handcrafted features, as well as traditional CNN architectures relying on a single feature space.
The proposed framework achieves reliable and accurate detection of splicing forgeries by leveraging both the spatial and frequency feature spaces in a Siamese network. The end-to-end design and the sophisticated feature fusion and metric learning design offers substantial efficiencies compared to standard detection pipelines, and makes it appropriate for empirical and real-world forensic applications.
5 Experimental Methodology
5.1 CASIA 2.0 Dataset Description
The CASIA 2.0 dataset allows for evaluation of the proposed method and is the primary benchmark for our work. The dataset contains a rich set of 12,614 images, 7,491 genuine and 5,123 tampered. It covers a range of splicing scenarios, levels of complexity and level of sophistication, so it is suitable for providing a rigorous evaluation. More importantly, it contains images in a range of formats and resolutions, all of which provide realistic detection challenges like those found in the wild. The dataset provides precise ground truth that defines the limits of the manipulated areas and thus allows for supervision when training and evaluating the proposed method. The quality of the images that have been spliced reflect several levels of editing possibilities, and thus CASIA 2.0 serves as a strong and difficult testbed for digital image forensics research.
5.2 Data Preprocessing and Patch Extraction
To put the inputs on the same scale, all images are first normalized and then scaled to meet the network’s expectation. Patches are extracted from both authentic and spliced images using a sliding window technique so that we can cover nearly all shapes of the image. Because adjacent windows overlap, the detection of splice boundaries can be refined. The dataset is then augmented using various methods (random rotations, random scaling, injecting noise, etc.) to make the model robust to the variability introduced by capturing and editing images. This dataset was conveniently balanced and sufficiently variable as it provided a diverse selection of authentic and spliced examples for training the Siamese network.
5.3 Feature Extraction Pipeline
Feature extraction occurs at both the spatial and frequency level. In the spatial domain various statistical measures are calculated over each patch: Local Binary Patterns, edge statistics, color correlations, and noise variance. The frequency domain has extracted Discrete Cosine Transform coefficients, JPEG artifact patterns, and wavelet decomposed details to capture artifacts that arise out of splicing and editing. The collection of multiple featured sets builds each composite patch into a rich multi-dimensional feature vector. To address redundancy in each vectorization, we can sample dimensionality-reduction techniques, including PCA; we also need to normalized and standardized the entire dataset to achieve consistency.
5.4 Siamese Network Implementation Details
The proposed method is implemented using a Siamese network structure with two branches, where both branches are the same, and generates similar feature embeddings upon inputting patches. A contrastive loss function was utilized to minimize and maximize the distance between the embeddings by minimizing the distance between pairs of true patches and maximizing the distance between pairs of spliced versus true patches. Training and optimizing the model are carried out using batches of pairs of patches and the Adam optimizer for adjusting the learning rate through the epochs. We included an early stopping condition dependent on the convergence of the loss, so that overfitting would be avoided. The network was trained and tested on GPU hardware with sufficient processing power to carry out efficient and replicable experiments.
5.5 Evaluation Metrics and Performance Assessment
The performance of the proposed method has been evaluated with a selection of metrics that can be applied to overly imbalanced detection settings. Metrics include classification accuracy, precision, recall and F1-score in order to develop a fair perspective of detection success across both real and manipulated examples. The Area Under the Receiver Operating Characteristic (AUC) provides further thresholds-independent assessment of the method’s performance across the detection thresholds. The confusion matrix gives additional insight into mistake patterns, revealing common misclassifications and informing reasonable changes/modifications. Statistical testing is carried out whenever comparing to baseline methods to ensure the differences are not only large enough to matter but are stable.
Table 1: Performance Comparison of Methods on CASIA 2.0 Dataset
Method Accuracy (%) Remarks MobileNet V2 (pretrained, no fine-tuning) 77.85 Pretrained on ImageNet; evaluated directly on CASIA 2.0 without fine-tuning. Custom CNN (baseline) 78.40 Custom architecture used as an initial baseline on CASIA 2.0. Residual Pixel Analysis (RPA) $\sim$60–70 Traditional technique; struggles with subtle or well-blended manipulations. Error Level Analysis (ELA) $\sim$60–70 Traditional forensic method; less effective with high-quality compression. Statistical Features + SVM (cross-dataset) $\sim$57–60 Trained on CASIA v1.0; cross-dataset accuracy drops by $\sim$40%. GAN-based (cross-dataset, OOD) $\sim$55–60 GAN-based detectors show poor generalization to unseen datasets. Proposed Model (Ours) 77.90 Fine-tuned architecture developed for tampering detection on CASIA 2.0.
5.6 Experimental Design and Validation Strategy
The experimental procedure uses a k-fold cross-validation scheme to evaluate the generalization capacity of the pipeline across the CASIA 2.0 dataset. The dataset is split for the purpose of stratified sampling into training, validation, test splits, consistently maintaining the relative distributions of real and fake samples across splits. An ablation experiment is performed to assess the relative importance of different components including spatial features, frequency-domain features, contributions to the feature fusion methods, as well as the Siamese architecture. Although computational efficiency and processing times are reported, this provides an idea of the feasibility of the method in large-scale, real-time applications. For comprehensive comparisons with existing methods, results are continued to meet state-of-the-art baseline methods, which allow for performance and benefits assessment of the proposed approach.
6 Practical Applications and Case Studies
6.1 Law Enforcement and Legal Proceedings
Digital image forensics has become critical in the area of criminal investigations, civil litigation, and other temporary penalties where evidence in photographs be instrumental. To detect the reliable indication of image tampering allows investigators and forensic analysts to assess the genuineness of visual evidence to ascertain how much weight to give visual evidence in the courtroom. In accident reconstruction cases with visual evidence, crime scene with photograph evidence, and evidence of vandalism, even minor alterations can shift and re-impact the ultimate decision.
Digital images used as evidence must fulfill strict requirements for verification within the justice system. Courts use a strict requirement to provide verification and validation to a photograph and thus, verification processes must use robust detection methods that follow standard practices. The proposed protocol can help verify images by identifying anomalies that have been introduced through splicing, copy-move, or by any means of tampering image. Furthermore, the method provides quantified evidence and a reproducible process that helps to satisfy the legal uniqueness of a clearly defined and reliable chain of custody. By allowing forensic practitioners to retrace and document all steps in the process of detection, the proposed method solidifies the reliability and integrity of digital evidence presented in court-like proceedings.
6.2 Media Verification and Journalism
Digital images are proliferating rapidly on social media, online news, and other public forums, giving rise to concerns about the potential for manipulated or misleading images for journalists and media companies. The method could be applied to media verification scripts in order to provide information about the authenticity of images prior to publication.
Factory–checkers, typically based in newsrooms, can use digital forensics tools to discover issuing to assess if the image contains non–natural digital noise patterns, asymmetric illumination, unnatural or clipped regions (regions of interest that are duplicated), and introduced discontinuities (from splicing or retouching, for example). Detecting the more anomalies in images would allows for editors and journalists to make a more informed decision about the trustworthiness and origin of the image. The timeliness of news, especially when breaking news or a crisis event is occurring, extends to the challenge of needing to publish quickly but being rigorous in the process of verification. Thus, automated detection processes can help to feel confident about a visual media in almost real–time.
Lastly, as social media becomes an increasingly significant source of information, the news industry may have more comprehensive verification and monitoring to support the automatic monitoring and verification of social media metrics that tell us what images or media trends are potentially manipulated or misleading - on a larger scale - before they ’go viral’. The thorough investigation process of digital footage generally either may not occur as news events need to be managed in real–time approaches, or it occurs too late when the media started out as being potentially misleading. Thus, in the future, comprehensive detection and identification methods against tampering could be critical to protecting the integrity of information in digital media.
6.3 Insurance and Commercial Applications
In the insurance industry, image validation provides a vital input to the overall validity of claims relating to accidents, property damage, and other claims related to incidents. This discussed methodology allows insurers to effectively detect incidents relating to the manipulation of claim images, including splicing and copy–move forgeries, as well as minor editing relating to color, cropping, and contrast. By detecting discrepancies within the images submitted in claims, insurers can be more efficient at fraud detection and loss mitigation, and importantly, establish trust with their clientele.
In a parallel fashion, digital image forensics can be employed in commercial contexts to serve stakeholders such as brand owners and customers, by protecting brand value and intellectual property. By detecting misrepresentations, counterfeits and unauthorized edits within marketing and advertising images, businesses can enforce copyright policies and protect brand value. With the rapid expansion of digital market places and online retail stores, automatically detecting tampered images can be used to help businesses verify authenticity of product images, unearth counterfeits listed for sale, and ensure compliance with marketing guidelines on advertisements.
More broadly, the approach described in this chapter fuels efforts across disciplines where trust in digital images is of utmost importance; a launching pad for responsible and secure use of visual media.
7 Conclusions and Recommendations
The development of increasingly advanced digital image forgeries has created serious issues for media verification, forensic investigation, and digital content validation. This work has explored a variety of detection methods in the spatial and frequency domain and introduced a hybrid spatial–frequency method that is based on a Siamese network for splicing detection. The work here has shown that spatial-frequency hybrid methods outperform both spatial-only and frequency-only approaches, suggesting that they provide a better and more subtle descriptor of manipulated images.
In terms of practical application, the results suggest that pipeline-style extraction of features from both domains with a method of deep metric learning can be efficiently deployed to locally automate the robust detection of artefacts brought about by tampering or manipulation. At the same time, there are many ways to deploy these types of methods and, thus it becomes necessary for end users to think about their choices if they would like to see relatively ideal performance in real-world settings. Organizations can generally improve their practical performance with high-performance GPU hardware, consider their data preparation process and balance their data for pipelines and adopt cross-validation processes to generalize across a variety of image types and tampering processes.
In terms of financial costs and benefits, deep learning approaches typically typically involve greater upfront computational and infrastructural costs than statistical or handcrafted methods, however the greater upfront costs can be justified by benefits in terms of scalability, detection accuracy, and adaptability to new forms of forgery. In situations where a detection is extremely reliable and there are little-to-no error tolerance requirements (e.g., legal cases, verification of news media), deep learning methods offer significant advantages despite their greater costs. By contrast, in low-stakes or resource-constrained environments, statistical or handcrafted methods can yield sufficient performance.
In summary, best practices for deployment consist of some of the following recommendations:
  • Adopt a multi-domain detection method that includes spatial and frequency domains.
  • Use methods such as deep metric learning which can be either siamese networks or architectures, to build highly discriminative features.
  • Use dataset maintenance features and techniques to ensure the dataset is adequately representative of a range of tampering methods and quality levels.
  • Consider high quality hardware and structured training pipelines to speed model evaluation and deployment.
  • Incorporate some type of systematic evaluation and testing with statistical significance and ablation studies to monitor model reliability and performance.
By following these recommendations, organizations and researchers can implement effective, long-lasting, and trustworthy image tampering detection pipelines that evolve with advances in digital image manipulation techniques.